OpenStudy (anonymous):

What value of n would make this expression equal to 1/81? 9^0 x 9^-4 x 9^n A) 0 B) 1 C) 2 D) 3 With work or explanation please! :)

OpenStudy (jdoe0001):

\(\bf aa^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\qquad thus \\ \quad \\ \cfrac{1}{81}\implies \cfrac{1}{9^{\color{red}{ 2}}}\implies 9^{\color{red}{ -2}}\qquad thus \\ \quad \\ 9^0\cdot 9^4\cdot 9^n\implies 9^{0+4+n}\qquad which\ equals\ 9^{\color{red}{ -2}}\qquad thus \\ \quad \\ 9^{0+4+n}=9^{\color{red}{ -2}}\implies n=?\) what do you think?

OpenStudy (jdoe0001):

hmmm got a typo but anyhow \(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\qquad thus \\ \quad \\ \cfrac{1}{81}\implies \cfrac{1}{9^{\color{red}{ 2}}}\implies 9^{\color{red}{ -2}}\qquad thus \\ \quad \\ 9^0\cdot 9^4\cdot 9^n\implies 9^{0+4+n}\qquad which\ equals\ 9^{\color{red}{ -2}}\qquad thus \\ \quad \\ 9^{0+4+n}=9^{\color{red}{ -2}}\implies n=?\) so.... what do you think?